4-3x=16 is the answer 4
dont get this one -13(3x-11)+2=28

6-4(2-1/2x)+x=22
help me please

Answer by checkley71(8403)

4-3X=16
-3X=16-4
-3X=12
X-12/-3
X=-4
--------------------------------------
-13(3X-11)+2=28
-39X+143+2=28
-39X=28-2-143
-39X=-117
X=-117/-39
X=3
--------------------------------------------------------
6-4(2-1/2X)+X=22
6-8+2X+X=22
3X=22-6+8
3X=24
X=24/3
X=8